# Radio Map-Assisted CSI Tracking with Uncertain Locations in Massive MIMO Networks

### v1

@@
Abstract
@@ II-B Observation Model
Figure 1: A MIMO system with $Q$ BS s equipped with MIMO antennas. The users move along the roads, and sparse measurements are used for constructing a channel covariance map which is a mapping from locations ${\bf p}\in\mathbb{R}^{3}$ in the $i$th grid to a channel covariance matrix $\mathbf{C}_{i}^{(q)}\in\mathbb{R}^{N_{t}\times N_{t}}$.
@@ IV-C Reduced Pilots and Batch Processing for Channel Tracking
$\tilde{{\bf y}}_{t}^{(q)}=[({\bf y}_{t-\tilde{T}+1}^{(q)})^{\text{T}},({\bf y}_{t-\tilde{T}+2}^{(q)})^{\text{T}},\dots,({\bf y}_{t}^{(q)})^{\text{T}}]^{\text{T}}\in\mathbb{C}^{M_{\text{b}}\times 1}$ (32)
where $M_{\text{b}}$ is the dimension of the virtual observation. Even though $M\ll N_{t}$, one can always choose a proper batch size to exploit the channel distributions.
@@
$\tilde{{\bf A}}=[{\bf A}_{t-\tilde{T}+1}^{\text{T}},{\bf A}_{t-\tilde{T}^{\prime}+2}^{\text{T}},\dots,{\bf A}_{t}^{\text{T}}]^{\text{T}}\in\mathbb{C}^{M_{\text{b}}\times N_{t}}$ (33)
@@ V-A Environment Setup and Scenarios
By simulating random walk along roads in the map, the radio database is established and it contains $20269$ user locations with its related CSI information. The user trajectories are randomly generated on the street. The observation ${\bf y}_{t}^{(q)}$ at each location ${\bf p}_{t}$ is generated using (4) where the actual channel ${\bf h}_{t}^{(q)}$ is generated using the ray-tracing technique of Wireless Insite®. To depict the uncertainty of locations, a random location shift following a Gaussian distribution $\mathcal{N}(0,400)$ is added to the actual user location ${\bf p}_{t}$. The grid length of the radio map is $5$ meters, the user velocity is set as $14$ meters per second, and the sampling period of each observation is $50$ milliseconds.
Figure 2: (a) Simulation environment where there are $7$ BS s equipped with $32$ dual-dipole MIMO antennas. (b) MIMO antennas with $N_{t}=64$ antenna arrays.
@@
V-B Radio Map Construction, Trajectory Discovery and Channel Estimation
With the assistance of radio map $\mathcal{M}$, fig. and fig. illustrate the localization error and channel capacity efficiency ratio versus the number of observations, respectively. Specifically, fig. shows the localization error versus the number of observations under different user velocity. It is observed that …. Fig. shows the channel capacity efficiency ratio versus the number of observations for different schemes. It is found that ….
Fig. displays the channel capacity efficiency ratio versus the average SNR of the communication environment.
Fig. shows the channel capacity efficiency ratio versus the number of observations for constructing the radio map $\mathcal{M}$ in an LOS environment where all the observations are LOS to the BS s. Fig. shows the channel capacity efficiency ratio versus the number of observations for constructing the radio map $\mathcal{M}$ in a NLOS environment where all the observations are NLOS to the BS s.
Figure 3: Radio map-assisted trajectory discovery and channel estimation. (a) Channel capacity efficiency ratio versus SNR under $2$ LOS BS s; (b) Channel capacity efficiency ratio versus SNR under $3$ NLOS BS s; (c) Localization error versus number of observations and type of BS s.
Figure 4: Performance versus the number of historical observations used for initializing the radio map. (a) Localization error versus number of historical observations; (b) Channel capacity efficiency ratio versus number of historical observations; (c) Localization error versus sampling period of each observation.
@@ Appendix A Proof of Theorem 1
According to the observation model in (4), the PDF of $p({\bf y}_{T}|\tilde{{\bf h}})$ is given by
$p({\bf y}_{T}|\tilde{{\bf h}})=\frac{\text{exp}(-({\bf y}_{T}-{\bf A}\tilde{{\bf h}})^{\text{H}}({\bf y}_{T}-{\bf A}\tilde{{\bf h}})/(2\sigma_{\text{n}}^{2}))}{(2\pi)^{M/2}\sigma_{\text{n}}}.$ (34)
Therefore, the log-likelihood of $p({\bf y}_{T}|\tilde{{\bf h}})$ satisfies
$\log p({\bf y}_{T}|\tilde{{\bf h}})\propto-\frac{\|{\bf y}_{T}-{\bf A}\tilde{{\bf h}}\|_{2}^{2}}{\sigma_{\text{n}}^{2}}.$ (35)
In addition, since the channel ${\bf\tilde{{\bf h}}}$ follows a complex Gaussian distribution in (3), and the deterministic channel covariance ${\bf C}({\bf p}_{T})$ is only dependent on the user location ${\bf p}_{\text{T}}$, the PDF of $p(\tilde{{\bf h}}|{\bf p}_{T})$ is given by
$p(\tilde{{\bf h}}|{\bf p}_{T})=\frac{\text{exp}(-\tilde{{\bf h}}^{\text{H}}({\bf C}({\bf p}_{T})+\sigma_{\text{h}}^{2}\mathbf{I}_{N_{t}\times N_{t}})^{-1}\tilde{{\bf h}}/2)}{(2\pi)^{N_{t}/2}|{\bf C}({\bf p}_{T})+\sigma_{\text{h}}^{2}\mathbf{I}_{N_{t}\times N_{t}}|^{1/2}}.$ (36)
@@
$\log p(\tilde{{\bf h}}|{\bf p}_{T})\propto-\tilde{{\bf h}}^{\text{H}}({\bf C}({\bf p}_{T})+\sigma_{\text{h}}^{2}\mathbf{I}_{N_{t}\times N_{t}})^{-1}\tilde{{\bf h}}.$ (37)
Given (37) and (35), the log-likelihood of $p({\bf y}_{T}|{\bf p}_{T})$ satisfies that
$\log p({\bf y}_{T}|{\bf p}_{T})\\ \propto-\frac{\|{\bf y}_{T}-{\bf A}\tilde{{\bf h}}\|_{2}^{2}}{\sigma_{\text{n}}^{2}}-\tilde{{\bf h}}^{\text{H}}({\bf C}({\bf p}_{T})+\sigma_{\text{h}}^{2}\mathbf{I}_{N_{t}\times N_{t}})^{-1}\tilde{{\bf h}}.$ (38)
Therefore, maximizing $\log p({\bf y}_{\text{T}}|{\bf p}_{T})$ is equivalent to maximizing $-\|{\bf y}_{T}-{\bf A}{\bf h}\|_{2}^{2}/\sigma_{\text{n}}^{2}-{\bf h}^{\text{H}}({\bf C}_{\pi_{T}}+\sigma_{\text{h}}^{2}\mathbf{I}_{N_{t}\times N_{t}})^{-1}{\bf h}$. Consequently, the problem (16) is equivalently rewritten as
$\mathop{\mbox{mimimize}}\limits_{{\bf h}}$ $\quad\frac{\|{\bf y}_{T}-{\bf A}\tilde{{\bf h}}\|_{2}^{2}}{\sigma_{\text{n}}^{2}}+\tilde{{\bf h}}^{\text{H}}({\bf C}({\bf p}_{T})+\sigma_{\text{h}}^{2}\mathbf{I}_{N_{t}\times N_{t}})^{-1}\tilde{{\bf h}}$ (39)
@@
$H(\tilde{{\bf h}})\triangleq\frac{\|{\bf y}_{T}-{\bf A}\tilde{{\bf h}}\|_{2}^{2}}{\sigma_{\text{n}}^{2}}+\tilde{{\bf h}}^{\text{H}}({\bf C}({\bf p}_{T})+\sigma_{\text{h}}^{2}\mathbf{I}_{N_{t}\times N_{t}})^{-1}\tilde{{\bf h}}.$ (40)
The gradient of $H(\tilde{{\bf h}})$ is given by
$\nabla_{\tilde{{\bf h}}}H(\tilde{{\bf h}})=-2{\bf A}^{\text{H}}({\bf y}_{T}-{\bf A}\tilde{{\bf h}})/\sigma_{\text{n}}^{2}+2({\bf C}({\bf p}_{T})+\sigma_{\text{h}}^{2}\mathbf{I}_{N_{t}\times N_{t}})^{-1}\tilde{{\bf h}}.$ (41)
Setting $\nabla_{{\bf\tilde{{\bf h}}}}H(\tilde{{\bf h}})$ to 00, the globally optimal solution ${\bf\tilde{{\bf h}}}^{*}$ to problem (16) and (39) is given by
$\tilde{{\bf h}}^{*}=({\bf A}^{\text{H}}{\bf A}+\sigma_{\text{n}}^{2}({\bf C}({\bf p}_{T})+\sigma_{\text{h}}^{2}{\bf I}_{N_{t}\times N_{t}})^{-1})^{-1}{\bf A}^{\text{H}}{\bf y}_{T}.$ (42)
@@ Appendix B Proof of Proposition (1)
The eigenvalue decomposition of $\tilde{{\bf C}}_{i}$ is given by
$\tilde{{\bf C}}_{i}$ $={\bf C}_{i}+\sigma_{\text{h}}^{2}{\bf I}_{N_{t}\times N_{t}}$
$={\bf U}_{i}\bm{\Sigma}_{i}{\bf U}_{i}^{\text{H}}+\sigma_{h}^{2}{\bf U}_{i}{\bf U}_{i}^{\text{H}}$ (43)
$={\bf U}_{i}(\bm{\Sigma}_{i}+\sigma_{\text{h}}^{2}{\bf I}_{N_{t}\times N_{t}}){\bf U}_{i}^{\text{H}}.$
where $\bm{\Sigma}_{i}=\text{diag}\{\lambda_{1},\lambda_{2},\dots,\lambda_{M}\}$ in which, $\lambda_{m}$ is the eigenvalue of ${\bf C}_{i}$ with $\lambda_{1}\geq\lambda_{2}\geq\cdots\geq\lambda_{M}$, and ${\bf U}_{i}$ is the matrix consisting of corresponding eigenvectors.
Given the optimal solution ${\bf A}^{*}$ in (20), the variance of ${\bf y}_{i}$ is given by
$\mathbb{V}\{{\bf A}{\bf h}_{i}\}$ $={\bf A}\tilde{{\bf C}}_{i}{\bf A}^{\text{H}}$
$={\bf U}_{i}(\bm{\Sigma}_{i}+\sigma_{\text{h}}^{2}{\bf I}_{N_{t}\times N_{t}}){\bf U}_{i}^{\text{H}}$
$={\bf U}_{i,M}^{\text{H}}{\bf U}_{i}(\bm{\Sigma}_{i}+\sigma_{\text{h}}^{2}{\bf I}_{N_{t}\times N_{t}}){\bf U}_{i}^{\text{H}}{\bf U}_{i,M}$
=$\overset{(i)}{=}[{\bf I}_{M\times M},{\bf 0}_{M\times N_{t}}](\bm{\Sigma}_{i}+\sigma_{\text{h}}^{2}{\bf I}_{N_{t}\times N_{t}})[{\bf I}_{M\times M},{\bf 0}_{M\times N_{t}}]^{\text{H}}$ (44)
where =$\overset{(i)}{=}$ is because that ${\bf U}_{i}$ is an is an orthogonal matrix.
Therefore, $\mathbb{V}\{{\bf A}{\bf h}_{i}\}$ captures the largest $M$ eigenvalues of $\tilde{{\bf C}}_{i}$, and thus, is maximized.

### v2

@@
Radio Map-Assisted CSI Tracking with Uncertain Locations in Massive MIMO Networks
Yuanshuai Zheng and Junting Chen Junting Chen is the corresponding author. The work was supported in part by the Basic Research Project No. HZQB-KCZYZ-2021067 of Hetao Shenzhen-HK S&T Cooperation Zone, by the National Science Foundation of China under Grant No. 62171398 and No. 62293482, by the Shenzhen Science and Technology Program under Grant No. JCYJ20220530143804010, No. KQTD20200909114730003, and No. KJZD20230923115104009, by Guangdong Research Projects No. 2019QN01X895, No. 2017ZT07X152, and No. 2019CX01X104, by the Shenzhen Outstanding Talents Training Fund 202002, by the Guangdong Provincial Key Laboratory of Future Networks of Intelligence (Grant No. 2022B1212010001), by the National Key R&D Program of China with grant No. 2018YFB1800800, and by the Key Area R&D Program of Guangdong Province with grant No. 2018B030338001. School of Science and Engineering, Shenzhen Future Network of Intelligence Institute (FNii-Shenzhen), and
Guangdong Provincial Key Laboratory of Future Networks of Intelligence
The Chinese University of Hong Kong, Shenzhen, Guangdong 518172, P.R. China
?abstractname?
@@ II-B Observation Model
?figurename? 1: A MIMO system with $Q$ BS s equipped with MIMO antennas. The users move along the roads, and sparse measurements are used for constructing a channel covariance map which is a mapping from locations ${\bf p}\in\mathbb{R}^{3}$ in the $i$th grid to a channel covariance matrix $\mathbf{C}_{i}^{(q)}\in\mathbb{R}^{N_{t}\times N_{t}}$.
@@ III-D CSI Tracking with the Exploration of Channel Covariance Subspace
The probing matrix in (23) takes the advantages of both random exploration of channel and subspace exploitation of channel covariance. In addition, the observation number at each location is reduced to $1$. The technical details are shown in Algorithm (2). The constant $T_{0}$ controls the length of the random exploration period. In particular, the case $T_{0}=1$ corresponds to the scenario that the CSI tracking is performed from the second measurement whose probing matrix is partially generated based on the channel covariance at the first estimated location. In this case, the user localization performance is weakened because of insufficient exploration of the multipath components of the channel. On the other hand, the case $T_{0}=T$ implies that there is no CSI tracking stage, which fails to capture the main information of the channel, resulting in poor channel estimation. The numerical results in Section (V) confirms the above statements.
Input: Radio maps $\mathcal{M}^{(q)}$ for $q\in\mathcal{Q}$, a time slot $\delta_{T}$, variances $\sigma_{\text{h}}^{2}$ and $\sigma_{\text{n}}^{2}$, and an initial mobility probability $\mathbb{P}_{0}$.
  1. 1.
    Initialize a probability matrix $\mathbf{P}\in\mathbb{R}^{N\times T}$ and a pointer matrix ${\bf Q}\in\mathbb{R}^{N\times T}$.
  2. 2.
    Initialization at $t=1,2,\dots,T_{0}$:
    1. (a)
      Generate ${\bf A}_{1}^{(q)}\in\mathbb{C}^{1\times N_{t}}$ as a Gaussian random matrix, send pilots, and observe measurements ${\bf y}_{t}^{(q)}$.
    2. (b)
      Conduct Step 2 in Algorithm 1.
    3. (c)
      Let $\tau_{1}\leftarrow\text{argmax}_{i\in\{1,2,\dots,N\}}P_{i,1}$.7777
  3. 3.
    For $t=T_{0}+1,\dots,T$
    1. (a)
      Generate ${\bf A}_{t}^{(q)}=(\hat{{\bf u}}_{t,m}^{(q)})^{\text{H}}$ with probability $\mathbb{P}(\hat{{\bf u}}_{t,m}^{(q)}|\hat{{\bf C}}^{(q)}({\bf p}_{t}))$ in (22) where $\hat{{\bf C}}^{(q)}({\bf p}_{t})$ is computed in 21, and $\hat{{\bf u}}_{t,m}^{(q)}$ is the $m$th eigenvector of $\hat{{\bf C}}^{(q)}({\bf p}_{t})$.
    2. (b)
      Conduct Step 3(a)i in Algorithm 1.
    3. (c)
      Let $\tau_{t}\leftarrow\text{argmax}_{i\in\{1,2,\dots,N\}}P_{i,t}$.
  4. 4.
    Iteratively update $\tau_{t}\leftarrow Q_{\tau_{t+1},t+1}$ for $t=T-1,T-2,\dots,1$.
  5. 5.
    Output the estimated trajectory $\tilde{\mathcal{P}}_{T}=\{{\bf x}_{\tau_{1}},{\bf x}_{\tau_{2}},\dots,{\bf x}_{\tau_{T}}\}$, and the estimated channel $({\bf h}_{T}^{(q)})^{*}$ at $t=T$ using (17).
Algorithm 2 Single-dimension Observations Driven CSI Tracking with the Exploration of Channel Covariance Subspace
@@ IV-C Reduced Pilots and Batch Processing for Channel Tracking
$\tilde{{\bf y}}_{t}^{(q)}=[({\bf y}_{t-\tilde{T}+1}^{(q)})^{\text{T}},({\bf y}_{t-\tilde{T}+2}^{(q)})^{\text{T}},\dots,({\bf y}_{t}^{(q)})^{\text{T}}]^{\text{T}}\in\mathbb{C}^{\tilde{T}\times 1}$ (32)
where $\tilde{T}$ is the batch size of the virtual observation. Even though $M\ll N_{t}$, one can always choose a proper batch size to exploit the channel distributions.
@@
$\tilde{{\bf A}}=[{\bf A}_{t-\tilde{T}+1}^{\text{T}},{\bf A}_{t-\tilde{T}^{\prime}+2}^{\text{T}},\dots,{\bf A}_{t}^{\text{T}}]^{\text{T}}\in\mathbb{C}^{\tilde{T}\times N_{t}}$ (33)
@@ V-A Environment Setup and Scenarios
By simulating random walk along roads in the map, the radio database is established and it contains $20269$ user locations with its related CSI information. The user trajectories are randomly generated on the street. The observation ${\bf y}_{t}^{(q)}$ at each location ${\bf p}_{t}$ is generated using (4) where the actual channel ${\bf h}_{t}^{(q)}$ is generated using the ray-tracing technique of Wireless Insite®. To depict the uncertainty of locations, a random location shift following a Gaussian distribution $\mathcal{N}(0,400)$ is added to the actual user location ${\bf p}_{t}$. The grid length of the radio map is $5$ meters, the user velocity is set as $50$ kilometer per hour (km/h), and the sampling period of each observation is $50$ milliseconds. The batch size $\tilde{T}$ is set as $20$.
?figurename? 2: (a) Simulation environment where there are $7$ BS s equipped with $32$ dual-dipole MIMO antennas. (b) MIMO antennas with $N_{t}=64$ antenna arrays.
@@
V-B Radio Map-assisted Trajectory Discovery and Channel Estimation
?figurename? 3: Channel capacity efficiency ratio versus SNR under $3$ LOS BS s.
?figurename? 4: Channel capacity efficiency ratio versus SNR under $3$ NLOS BS s.
?figurename? 5: Localization error versus number of observations and type of BS s.
With the assistance of radio map $\mathcal{M}^{(q)}$, fig. 3 and fig. 4 illustrate channel capacity efficiency ratio versus SNR. The number of observations is $400$, which corresponds to about $280$ meters as the velocity is $50$ km/h and the observations are collected every $50$ ms. It is found that the if all the channel data are utilized to construct the radio map, the radio map-assisted approach can achieve over $94\%$ performance of the perfect channel when SNR is equal to or greater than $10$ dB. The proposed Algorithm 3 utilizes $1800$ historical observations to initialize a radio map, and then it also achieves over $90\%$ channel capacity of that of perfect CSI. The AR and the LSTM approaches shows poor performance because they assume that the channel can be accurately represented with historical channels as shown in (34), and the performance of these approaches depend on the accuracy of estimation of historical channels.
Additionally, it is observed from fig. 3 and fig. 4 that the proposed radio map-assisted approach can also estimate the NLOS channels in an actual city environment. Although there are lots of scatters and reflections due to complex topological features, the radio map-assisted approach can still achieve over $94\%$ performance of the perfect channel under SNR $=20$ dB. The results from that the radio map has captured the deterministic channel distributions in presence of blockage.
Fig, 5 shows the localization error versus number of observations. The user trajectories are randomly generated, and they are LOS to BS 1 while NLOS to BS 4. It is found that the radio-map assisted approach has a localization error of less than $5$ meters when $30$ observations are collected if the user trajectory is LOS to a BS. For user trajectories in NLOS conditions, the localization error can still be less than $30$ meters when $70$ observations are collected.
V-C Radio Map Construction, Trajectory Discovery and Channel Estimation
?figurename? 6: Localization error versus number of historical observations used for initializing the radio map.
?figurename? 7: Channel capacity efficiency ratio versus number of historical observations used for initializing the radio map.
?figurename? 8: Localization error versus sampling period of each observation.
Fig.6 shows the localization error versus the number of historical observations for initializing the radio map. It is found that more historical observations contribute to smaller localization error in the initial stage. Additionally, it is observed that more BS s perform better user localization. Fig.7 displays the channel capacity efficiency ratio versus the number of historical observations. User trajectories in LOS propagation conditions can obtain better channel estimation performance while it is more difficult to estimate MIMO channels in NLOS conditions. Fig.8 shows the localization error versus sampling period. It is found that the even the BS can collect observations with a higher rate, the sampling rate does not affect the localization performance too much when there are enough historical observations for radio map construction.
@@ Appendix A Proof of Theorem 1
According to the observation model in (4), the PDF of $p({\bf y}_{T}^{(q)}|{\bf h}_{T}^{(q)})$ is given by
$p({\bf y}_{T}^{(q)}|{\bf h}_{T}^{(q)})=\frac{\text{exp}(-({\bf y}_{T}^{(q)}-{\bf A}{\bf h}_{T}^{(q)})^{\text{H}}({\bf y}_{T}^{(q)}-{\bf A}{\bf h}_{T}^{(q)})/(2\sigma_{\text{n}}^{2}))}{(2\pi)^{M/2}\sigma_{\text{n}}}.$ (35)
Therefore, the log-likelihood of $p({\bf y}_{T}^{(q)}|{\bf h}_{T}^{(q)})$ satisfies
$\log p({\bf y}_{T}^{(q)}|{\bf h}_{T}^{(q)})\propto-\frac{\|{\bf y}_{T}^{(q)}-{\bf A}{\bf h}_{T}^{(q)}\|_{2}^{2}}{\sigma_{\text{n}}^{2}}.$ (36)
In addition, since the channel ${\bf h}_{T}^{(q)}$ follows a complex Gaussian distribution in (3), and the deterministic channel covariance ${\bf C}^{(q)}({\bf p}_{T})$ is only dependent on the user location ${\bf p}_{T}$, the PDF of $p({\bf h}_{T}^{(q)}|{\bf p}_{T})$ is given by
$p({\bf h}_{T}^{(q)}|{\bf p}_{T})=\frac{\text{exp}(-({\bf h}_{T}^{(q)})^{\text{H}}({\bf C}^{(q)}({\bf p}_{T})+\sigma_{\text{h}}^{2}\mathbf{I}_{N_{t}\times N_{t}})^{-1}{\bf h}_{T}^{(q)}/2)}{(2\pi)^{N_{t}/2}|{\bf C}({\bf p}_{T})+\sigma_{\text{h}}^{2}\mathbf{I}_{N_{t}\times N_{t}}|^{1/2}}.$ (37)
@@
$\log p({\bf h}_{T}^{(q)}|{\bf p}_{T})\propto-({\bf h}_{T}^{(q)})^{\text{H}}({\bf C}^{(q)}({\bf p}_{T})+\sigma_{\text{h}}^{2}\mathbf{I}_{N_{t}\times N_{t}})^{-1}{\bf h}_{T}^{(q)}.$ (38)
Given (38) and (36), the log-likelihood of $p({\bf y}_{T}^{(q)}|{\bf p}_{T})$ satisfies that
$\log p({\bf y}_{T}^{(q)}|{\bf p}_{T})\\ \propto-\frac{\|{\bf y}_{T}^{(q)}-{\bf A}{\bf h}_{T}^{(q)}\|_{2}^{2}}{\sigma_{\text{n}}^{2}}-({\bf h}_{T}^{(q)})^{\text{H}}({\bf C}^{(q)}({\bf p}_{T})+\sigma_{\text{h}}^{2}\mathbf{I}_{N_{t}\times N_{t}})^{-1}{\bf h}_{T}^{(q)}.$ (39)
Therefore, maximizing $\log p({\bf y}_{T}^{(q)}|{\bf p}_{T})$ is equivalent to maximizing $-\|{\bf y}_{T}^{(q)}-{\bf A}{\bf h}_{T}^{(q)}\|_{2}^{2}/\sigma_{\text{n}}^{2}-({\bf h}_{T}^{(q)})^{\text{H}}({\bf C}^{(q)}({\bf p}_{T})+\sigma_{\text{h}}^{2}\mathbf{I}_{N_{t}\times N_{t}})^{-1}{\bf h}_{T}^{(q)}$. Consequently, the problem (16) is equivalently rewritten as
$\mathop{\mbox{mimimize}}\limits_{{\bf h}_{T}^{(q)}}$ $\quad\frac{\|{\bf y}_{T}^{(q)}-{\bf A}{\bf h}_{T}^{(q)}\|_{2}^{2}}{\sigma_{\text{n}}^{2}}$ (40)
$\quad+({\bf h}_{T}^{(q)})^{\text{H}}({\bf C}^{(q)}({\bf p}_{T})+\sigma_{\text{h}}^{2}\mathbf{I}_{N_{t}\times N_{t}})^{-1}{\bf h}_{T}^{(q)}$
@@
$H({\bf h}_{T}^{(q)})$ $\triangleq\frac{\|{\bf y}_{T}^{(q)}-{\bf A}{\bf h}_{T}^{(q)}\|_{2}^{2}}{\sigma_{\text{n}}^{2}}$ (41)
$+({\bf h}_{T}^{(q)})^{\text{H}}({\bf C}^{(q)}({\bf p}_{T})+\sigma_{\text{h}}^{2}\mathbf{I}_{N_{t}\times N_{t}})^{-1}{\bf h}_{T}^{(q)}.$
The gradient of $H({\bf h}_{T}^{(q)})$ is given by
$\nabla_{{\bf h}_{T}^{(q)}}H({\bf h}_{T}^{(q)})$ $=-2{\bf A}^{\text{H}}({\bf y}_{T}^{(q)}-{\bf A}{\bf h}_{T}^{(q)})/\sigma_{\text{n}}^{2}$ (42)
$+2({\bf C}^{(q)}({\bf p}_{T})+\sigma_{\text{h}}^{2}\mathbf{I}_{N_{t}\times N_{t}})^{-1}{\bf h}_{T}^{(q)}.$
Setting $\nabla_{{\bf h}_{T}^{(q)}}H({\bf h}_{T}^{(q)})$ to 00, the globally optimal solution ${\bf h}_{T}^{(q)}$ to problem (16) and (40) is given by
@@ Appendix B Proof of Proposition (1)
The eigenvalue decomposition of $\tilde{{\bf C}}_{i}^{(q)}={\bf C}_{i}^{(q)}+\sigma_{\text{h}}^{2}{\bf I}_{N_{t}\times N_{t}}$ is given by
$\tilde{{\bf C}}_{i}^{(q)}$ $={\bf C}_{i}^{(q)}+\sigma_{\text{h}}^{2}{\bf I}_{N_{t}\times N_{t}}$
$={\bf U}_{i}^{(q)}\bm{\Sigma}_{i}^{(q)}({\bf U}_{i}^{(q)})^{\text{H}}+\sigma_{\text{h}}^{2}{\bf U}_{i}^{(q)}({\bf U}_{i}^{(q)})^{\text{H}}$
$={\bf U}_{i}^{(q)}(\bm{\Sigma}_{i}^{(q)}+\sigma_{\text{h}}^{2}{\bf I}_{N_{t}\times N_{t}})({\bf U}_{i}^{(q)})^{\text{H}}.$ (44)
where $\bm{\Sigma}_{i}^{(q)}=\text{diag}\{\lambda_{i,1}^{(q)},\lambda_{i,2}^{(q)},\dots,\lambda_{i,M}^{(q)}\}$ in which, $\lambda_{i,m}^{(q)}$ is the eigenvalue of ${\bf C}_{i}^{(q)}$ with $\lambda_{i,1}^{(q)}\geq\lambda_{i,2}^{(q)}\geq\cdots\geq\lambda_{i,M}^{(q)}$, and ${\bf U}_{i}^{(q)}$ is the matrix consisting of corresponding eigenvectors. Define ${\bf U}_{i,M}^{(q)}$ as the collection of the first $M$ eigenvector of ${\bf C}_{i}^{(q)}$, i.e., ${\bf U}_{i,M}^{(q)}\triangleq[{\bf u}_{t,1}^{(q)},{\bf{\bf u}}_{t,2}^{(q)},\dots,{\bf u}_{t,M}^{(q)}]^{\text{H}}$.
Given the optimal solution ${\bf A}_{t}^{(q)}$ in (20), the variance of ${\bf y}_{t}^{(q)}$ with ${\bf p}_{t}\in\mathcal{G}_{i}$ is given by
$\mathbb{V}\{{\bf A}_{t}^{(q)}{\bf h}_{t}^{(q)}\}$ $={\bf A}_{t}^{(q)}\tilde{{\bf C}}_{i}^{(q)}({\bf A}_{t}^{(q)})^{\text{H}}$
$=({\bf U}_{i,M}^{(q)})^{\text{H}}{\bf U}_{i}^{(q)}(\bm{\Sigma}_{i}^{(q)}+\sigma_{\text{h}}^{2}{\bf I}_{N_{t}\times N_{t}})({\bf U}_{i}^{(q)})^{\text{H}}{\bf U}_{i,M}^{(q)}$
=$\overset{(i)}{=}\tilde{{\bf I}}_{M\times N_{t}}(\bm{\Sigma}_{i}^{(q)}+\sigma_{\text{h}}^{2}{\bf I}_{N_{t}\times N_{t}})\tilde{{\bf I}}_{M\times N_{t}}^{\text{H}}$ (45)
where =$\overset{(i)}{=}$ is because that ${\bf U}_{i}^{(q)}$ is an is an orthogonal matrix and $\tilde{{\bf I}}_{M\times N_{t}}=[{\bf I}_{M\times M},{\bf 0}_{M\times N_{t}}]$.
Therefore, $\mathbb{V}\{{\bf A}_{t}^{(q)}{\bf h}_{t}^{(q)}\}$ captures the largest $M$ eigenvalues of $\tilde{{\bf C}}_{i}^{(q)}$, and thus, is maximized.